# Resilient Topology-Aware Coordination for Dynamic 3D UAV Networks under Node Failure

### v1

@@ Abstract.
In 3D Aerial-Ground Integrated Networks (AGINs), ensuring continuous service coverage under unexpected hardware failures is critical for mission-critical applications. While Multi-Agent Reinforcement Learning (MARL) has shown promise in autonomous coordination, its resilience under sudden node failures remains a challenge due to dynamic topology deformation. This paper proposes a Topology-Aware Graph MAPPO (TAG-MAPPO) framework designed to enhance system survivability through autonomous 3D spatial reconfiguration. Our framework incorporates graph-based feature aggregation with a residual ego-state fusion mechanism to capture intricate inter-agent dependencies. This architecture enables the surviving swarm to rapidly adapt its topology compared to conventional Multi-Layer Perceptron (MLP) based approaches. Extensive simulations across heterogeneous environments, ranging from interference-limited Crowded Urban to sparse Rural areas, validate the proposed approach. The results demonstrate that TAG-MAPPO consistently outperforms baselines in both stability and efficiency; specifically, it reduces redundant handoffs by up to 50 percent while maintaining a lead in energy efficiency. Most notably, the framework exhibits exceptional self-healing capabilities following a catastrophic node failure. TAG-MAPPO restores over 90 percent of the pre-failure service coverage within 15 time steps, exhibiting a significantly faster V-shaped recovery trajectory than MLP baselines. Furthermore, in dense urban scenarios, the framework achieves a post-failure Jain’s Fairness Index that even surpasses its original four-UAV configuration by effectively resolving service overlaps. These findings suggest that topology-aware coordination is essential for the realization of resilient 6G aerial networks and provides a robust foundation for adaptive deployments in volatile environments.
@@
Aerial-Ground Integrated Networks (AGINs), Graph Attention Network, Network Resilience, Node Failure Recovery, Handoff Optimization
copyright: acmlicensedjournal: TOITjournalvolume: XXjournalnumber: XXarticle: XXjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Networks Mobile networksccs: Networks Wireless access points, base stations and infrastructureccs: Networks Network resources allocationccs: Computing methodologies Multi-agent reinforcement learning
1. Introduction
@@
The dawn of 6G networks signals a fundamental paradigm shift from terrestrial-dependent infrastructure to an integrated, three-dimensional Non-Terrestrial Network (NTN) architecture (Saad et al., 2020; Giordani and Zorzi, 2021). Within this aerial frontier, Unmanned Aerial Vehicles (UAVs) serving as Aerial Base Stations (ABSs) have emerged as a cornerstone solution due to their superior maneuverability, rapid on-demand deployment, and inherent ability to establish Line-of-Sight (LoS) connectivity (Mozaffari et al., 2019; Al-Hourani et al., 2014). These UAV-assisted networks are particularly indispensable in mission-critical scenarios where terrestrial infrastructure is compromised, such as disaster relief operations or large-scale urban events (Merwaday and Guvenc, 2015; Zhao et al., 2019). However, beyond the pervasive challenge of non-stationary user distributions, ensuring system resilience in the face of unexpected hardware malfunctions or catastrophic node failures remains a formidable obstacle for next-generation aerial orchestration.
@@
To sustain optimal coverage amidst environmental volatility, Multi-Agent Proximal Policy Optimization (MAPPO) has been widely adopted as a robust control framework (Yu et al., 2022; Cui et al., 2020). Nevertheless, a persistent limitation in reinforcement learning (RL) based UAV control is the structural rigidity of standard Multi-Layer Perceptron (MLP) architectures. Building upon our established research in MARL-based UAV control, our prior work (Lai, 2026) addressed the plasticity-stability dilemma by introducing a Group-Decoupled MAPPO (G-MAPPO) algorithm. While G-MAPPO effectively mitigated gradient conflicts across localized geographic regions, its efficacy was constrained by the assumption of a quasi-static network topology. Specifically, it assumed a constant number of active UAV nodes, leaving the system vulnerable to the topological disruptions inherent in real-world deployments. This paper extends that trajectory by pivoting from gradient decoupling to topology-aware coordination, specifically targeting the resilience gaps left by static structural assumptions.
@@
Practical aerial networks, however, are subject to dual-layer dynamics consisting of the continuous micro-temporal displacement of ground users and the discrete, catastrophic disruptions caused by sudden UAV node failures. MLP-based agents struggle with these topological ruptures due to their fixed input dimensions and inherent lack of permutation invariance. When a neighboring node fails, MLP architectures often suffer from coordination collapse because they cannot dynamically re-weight remaining spatial features, leading to service outages and a failure to re-establish coverage equilibrium.
@@
To overcome these structural limitations and achieve self-healing communication, this paper proposes a novel Topology-Aware Graph Attention (TAG) framework. Recent literature highlights that graph-based learning is uniquely suited for wireless resource management due to its inherent ability to generalize to varying numbers of nodes (Dai et al., 2025). Instead of perceiving the environment as a static feature vector, we model the multi-UAV system as a dynamic graph (Scarselli et al., 2009). By incorporating a sophisticated graph attention mechanism inspired by advancements in Graph Attention Networks (GAT) (Veličković et al., 2018; Vaswani et al., 2017), our framework empowers agents to dynamically assign importance weights to neighboring entities. This architecture, enhanced by random observation shuffling during training, enables the system to proactively adapt to both smooth user movements and abrupt topological changes such as the sudden loss of a coordinated agent.
@@
  • Topology-Aware Coordination Architecture: We propose a novel MARL framework that utilizes graph-based feature aggregation to capture the underlying relational structures of UAV networks. This architecture enables agents to perform spatial reasoning beyond simple numerical observations, leading to more stable coordination in interference-limited regimes.
@@
  • Signaling Stability and Resource Efficiency: Through the integration of a handoff-penalized reward function and energy-efficient policy learning, our framework achieves a superior balance between coverage expansion and signaling stability. Experimental data show that TAG-MAPPO reduces redundant handoffs by nearly 50% in sparse topologies while maintaining optimized propulsion and communication power consumption.
  • Autonomous Resilience and Self-Healing: We introduce a robust training paradigm incorporating Random Observation Shuffling (ROS), which empowers the network to adapt to dynamic agent populations. Our results demonstrate a unique self-healing capability where the system restores over 90% of service coverage following a random node failure.
@@
  • Fairness Optimization via Structural Reconfiguration: We provide a critical insight into the system’s ability to resolve service overlaps under resource-constrained conditions. Our analysis reveals that in crowded urban environments, the framework can strategically re-optimize its spatial deployment after a node failure to achieve a more balanced resource distribution, effectively surpassing its pre-failure fairness performance.
@@
The remaining sections of this paper are organized as follows. Section 2 reviews related work on UAV deployment and MARL for UAV swarms. Sections 3 and 4 respectively present the system model and problem formulation. Section 5 details the proposed TAG-MAPPO framework. Section 6 discusses the simulation setup and performance evaluation. Finally, Section 7 concludes the paper.
2. Related Work
@@
This section reviews the existing literature across three critical domains: UAV coverage optimization, network resilience under node failure, and the integration of graph-based attention mechanisms within multi-agent reinforcement learning (MARL).
2.1. Multi-UAV Control and Coverage Optimization
@@
The deployment of UAVs for wireless coverage has been extensively studied through heuristic, mathematical, and learning-based approaches. Early research typically formulated the coverage problem as a facility location task. Within this context, optimal 3D locations for static users were derived in (Mozaffari et al., 2016; Alzenad et al., 2017) by utilizing convex optimization and circle packing theory. To address energy constraints, trajectory optimization frameworks were proposed in (Zeng and Zhang, 2017) to balance throughput against propulsion energy consumption. More recently, meta-heuristic algorithms have been applied to dynamic scenarios. An evolving collaborative differential evolution algorithm was proposed in (Xu et al., 2025) for dynamic path planning, and a minimum-UAV coverage problem was addressed in (Gong et al., 2024) through energy-efficient graph algorithms. However, these model-based methods typically assume global Channel State Information (CSI) availability and suffer from high computational latency, which makes them less suitable for real-time control in volatile environments.
@@
With the advancement of deep reinforcement learning (DRL), decentralized control strategies have emerged as a promising alternative. Since the introduction of Multi-Agent Deep Deterministic Policy Gradient (MADDPG) (Lowe et al., 2017), various DRL-based controllers have been developed for aerial networks. DRL was utilized in (Hoang et al., 2025) for adaptive 3D placement in 6G airborne small cells, while a joint 3D trajectory design and resource scheduling framework was proposed in (Liu et al., 2024) for space-air-ground integrated networks. Beyond throughput maximization, ensuring fairness has gained significant attention. An adaptive deployment approach was proposed in (Lai et al., 2023) to balance offload traffic, highlighting the importance of fairness-aware protocols. Furthermore, a utility-driven route adaptation scheme was introduced in (Aryendu et al., 2025) for green cooperative networks. Despite these advancements, most existing works rely on coordinate-based Multi-Layer Perceptrons (MLPs). This architecture struggles to capture the permutation-invariant nature of UAV swarms (Zaheer et al., 2017), which often leads to coordination collapse when the number of neighbors fluctuates.
2.2. Network Resilience and Fault Tolerance
@@
In mission-critical scenarios such as disaster relief, system resilience is paramount. This refers specifically to the ability of a network to recover from sudden disruptions. Robustness in UAV search path planning has been enhanced using DRL for complex disaster environments (Liang et al., 2026). Additionally, a two-stage optimization approach was proposed in (Zheng et al., 2025) for collaborative post-disaster communications. A recent study in (Huang et al., 2026) addressed resilient topology configuration in 3D UAV networks by formulating the problem as a mixed-integer non-linear program (MINLP) and employing recursive optimization to maintain connectivity. While such model-based approaches provide theoretical performance bounds, they often entail significant computational overhead when handling large-scale topological shifts.
To adapt to such non-stationary environments, continual learning (CL) and distributed techniques have been adopted. More recently, a resilient topology optimization framework was introduced in (Zhang et al., 2026), which leverages quantum annealing to generate a diverse set of candidate topologies offline for rapid online switching. While effective in utilizing quantum parallelism for structural diversity, such approaches rely on pre-defined candidate sets and specialized hardware, which may limit their adaptability in unforeseen disaster scenarios where node status changes dynamically.
@@
Building on foundational CL algorithms, a spatiotemporal-aware DRL framework was introduced in (Chen et al., 2026) for cooperative coverage. In our prior work (Lai, 2026), a spatiotemporal continual learning framework based on Group-Decoupled MAPPO (G-MAPPO) was developed to enable UAVs to adapt to varying geographic user distributions. However, current research predominantly focuses on macro-spatial variations. Micro-temporal resilience involves the ability of a swarm to self-heal and reconfigure its topology immediately after a sudden node failure without relying on pre-computed candidate sets or recursive re-optimization. This domain remains under-explored, and this paper addresses the gap by proposing a topology-aware controller capable of real-time fault recovery through decentralized coordination.
2.3. Attention Mechanisms and Reconfigurable MARL
@@
To overcome the limitations of fixed-input architectures, Graph Neural Networks (GNNs) and attention mechanisms have been integrated into MARL. Recent reviews emphasize that graph-based learning is uniquely suited for wireless resource management due to its inherent ability to generalize to dynamic network statuses and scale with varying numbers of nodes (Dai et al., 2025). The effectiveness of using graph representations for resilient coordination is further supported by the topology sampling logic explored in (Zhang et al., 2026). The Transformer architecture (Vaswani et al., 2017) demonstrated the power of attention in sequence modeling, and it has been further shown that graph-based approaches possess strong relational inductive biases suitable for physical systems (Battaglia et al., 2018).
@@
In the context of UAV mobility management, a hierarchical multi-agent DRL approach was proposed in (Meer et al., 2026) to handle dynamic clustering. Although hierarchical structures can be effective, they are often complex to train. Attention mechanisms offer a more scalable solution. Algorithms such as MAAC (Iqbal and Sha, 2019) and MAGIC (Niu et al., 2021) use attention to learn dynamic communication graphs. Nevertheless, few existing works explicitly integrate relative velocity into the attention mechanism for resilient trajectory control. By embedding velocity vectors and spatial relations, the proposed TAG-MAPPO framework enables agents to perceive the evolution of the network topology. This approach enhances stability in high-mobility environments and ensures that the remaining swarm can autonomously re-optimize coverage after a catastrophic node failure.
3. System Model
3.1. 3D Aerial-Ground Network Architecture
3.1.1. Scenario Description
@@
We consider a dynamic High-Mobility Edge Coverage scenario within 6G Non-Terrestrial Networks (NTN). As illustrated in Figure 1, a fleet of Unmanned Aerial Vehicles (UAVs) operates as a flying small-cell tier to assist a macro Ground Base Station (GBS). This architecture is specifically designed for dynamic clusters such as vehicular platoons or emergency fleets, effectively addressing severe spatiotemporal non-stationarity. The UAV swarm dynamically tracks user mobility to offload traffic and maintain reliable Line-of-Sight (LoS) connectivity, thereby mitigating the inherent coverage limitations of terrestrial infrastructure.
Fig 1. Illustration of the dynamic 3D aerial-ground integrated network scenario. A fleet of UAVs functions as flying base stations to assist the Macro GBS in serving high-mobility vehicular platoons at a traffic intersection. The red solid arrows represent the high-capacity wireless backhaul links connected to the GBS (Data Plane), forming a star topology. The blue dotted lines denote the inter-UAV coordination links for exchanging local state information (Control Plane). The dashed ovals indicate the directional service footprints dynamically tracking the moving user clusters.
3.1.2. Physical Configuration and Survivability
@@
The system comprises $K_{U}$ rotary-wing UAVs and one stationary Ground Base Station (GBS), totaling $K=K_{U}+1$ serving nodes. These nodes are deployed to serve $M$ ground users distributed over a square target area $\mathcal{D}\subset\mathbb{R}^{2}$. Let $\mathcal{K}_{U}=\{1,\dots,K_{U}\}$ denote the set of mobile UAVs and $\mathcal{M}=\{1,\dots,M\}$ denote the set of users. The GBS is situated at the center of the target area $\mathbf{p}_{\mathrm{GBS}}=[x_{c},y_{c},0]^{T}$ to provide stable backhaul connectivity.
@@
To characterize system resilience, we define a node survival indicator $\xi_{k}(t)\in\{0,1\}$ for each UAV $k\in\mathcal{K}_{U}$. If UAV $k$ is operational at time $t$, $\xi_{k}(t)=1$; otherwise, $\xi_{k}(t)=0$ represents a sudden hardware failure or energy depletion. Since the GBS is stationary and assumed to have a persistent power supply, we define $\xi_{\mathrm{GBS}}(t)=1$ for all $t$.
@@
Based on these indicators, the active subset of serving nodes at time $t$ is defined as $\mathcal{V}(t)=\{k\in\mathcal{K}_{U}\mid\xi_{k}(t)=1\}\cup\{\mathrm{GBS}\}$. An inactive UAV ($\xi_{k}(t)=0$) is characterized by the cessation of all locomotion and transmission functions, effectively being removed from the coordination process and the interference environment. The state of each operational UAV $k\in\mathcal{V}(t)\setminus\{\mathrm{GBS}\}$ is defined by its 3D coordinates $\mathbf{p}_{k}(t)=[x_{k}(t),y_{k}(t),z_{k}(t)]^{T}$ and its instantaneous velocity vector $\mathbf{v}_{k}(t)$, which is subject to a maximum flight speed constraint $\|\mathbf{v}_{k}(t)\|\leq V_{\max}^{\text{UAV}}$. Furthermore, UAV flight altitudes are constrained within a safe operational corridor $H_{\min},H_{\max}]$.
3.1.3. Inter-UAV Coordination and Control Plane

### v2

@@ Abstract
Ensuring continuous service coverage under unexpected hardware failures is a fundamental challenge for 3D Aerial-Ground Integrated Networks. Although Multi-Agent Reinforcement Learning facilitates autonomous coordination, traditional architectures often lack resilience to sudden topology deformations. This paper proposes the Topology-Aware Graph MAPPO (TAG-MAPPO) framework to enhance system survivability through autonomous 3D spatial reconfiguration. Our framework integrates graph-based feature aggregation with a residual ego-state fusion mechanism to capture intricate inter-agent dependencies. To achieve structural robustness, we introduce a Random Observation Shuffling mechanism that fosters strong generalization to agent population fluctuations by breaking coordinate-index dependencies. Extensive simulations across heterogeneous environments, including high-speed mobility at 15 meters per second, demonstrate that TAG-MAPPO significantly outperforms Multi-Layer Perceptron baselines. Specifically, the framework reduces redundant handoffs by up to 50 percent while maintaining superior energy efficiency. Most notably, TAG-MAPPO exhibits exceptional self-healing capabilities, restoring over 90 percent of pre-failure coverage within 15 time steps. In dense urban scenarios, the framework achieves a post-failure fairness index surpassing its original four-UAV configuration by autonomously resolving service overlaps and interference. These findings confirm that topology-aware coordination is essential for resilient 6G aerial networks.
Index Terms:
@@
Aerial-Ground Integrated Networks (AGINs), Graph Attention Network, Network Resilience, Node Failure Recovery, Handoff Optimization, Energy Efficiency, Fairness, Multi-Agent Reinforcement Learning (MARL).
I Introduction
@@
The dawn of 6G networks signals a paradigm shift from terrestrial-dependent infrastructure to integrated, three-dimensional Non-Terrestrial Network (NTN) architectures [23, 8]. Within this frontier, Unmanned Aerial Vehicles (UAVs) serving as Aerial Base Stations (ABSs) are a cornerstone solution due to their maneuverability, rapid deployment, and Line-of-Sight (LoS) connectivity [21, 1]. These networks are indispensable in mission-critical scenarios where terrestrial infrastructure is compromised, such as disaster relief or large-scale urban events [19, 33]. However, beyond non-stationary user distributions, ensuring system resilience against hardware malfunctions or catastrophic node failures remains a formidable obstacle for aerial orchestration.
@@
To sustain coverage amidst volatility, Multi-Agent Proximal Policy Optimization (MAPPO) has been widely adopted [28, 6]. Nevertheless, a persistent limitation in RL-based UAV control is the structural rigidity of Multi-Layer Perceptron (MLP) architectures. Building upon our prior research, our previous work [14] addressed the plasticity-stability dilemma via Group-Decoupled MAPPO (G-MAPPO). While G-MAPPO mitigated gradient conflicts, its efficacy was constrained by the assumption of a quasi-static topology with a constant number of active nodes, leaving systems vulnerable to real-world disruptions. This paper extends that trajectory by pivoting toward topology-aware coordination to target resilience gaps left by static structural assumptions.
@@
Practical aerial networks face dual-layer dynamics: continuous micro-temporal user displacement and discrete, catastrophic disruptions from UAV failures. MLP-based agents struggle with these ruptures due to fixed input dimensions and a lack of permutation invariance. When a neighbor fails, MLP architectures often suffer coordination collapse as they cannot dynamically re-weight spatial features, leading to service outages and failure to re-establish equilibrium.
@@
To achieve self-healing communication, this paper proposes the Topology-Aware Graph Attention (TAG) framework. Graph-based learning is uniquely suited for wireless management due to its ability to generalize to varying node counts [7]. Instead of static feature vectors, we model the multi-UAV system as a dynamic graph [24]. By incorporating a graph attention mechanism inspired by Graph Attention Networks (GAT) [26, 25], agents can dynamically assign importance weights to neighbors. This architecture, enhanced by a Random Observation Shuffling (ROS) mechanism, enables the system to adapt to both user movements and abrupt topological changes.
@@
  • Topology-Aware Coordination Architecture: We propose a MARL framework utilizing graph-based feature aggregation to capture relational structures. This enables spatial reasoning beyond simple observations, leading to stable coordination in interference-limited regimes.
@@
  • Signaling Stability and Resource Efficiency: Integrating a handoff-penalized reward function with energy-efficient policy learning achieves a superior balance between coverage and signaling stability. Results show TAG-MAPPO reduces redundant handoffs by nearly 50 percent in sparse topologies while maintaining optimized energy efficiency.
  • Autonomous Resilience via ROS: We introduce a robust training paradigm incorporating ROS, regularizing the policy to be permutation-invariant and resilient to population fluctuations. By internalizing coordination logic under stochastic failures, the framework generalizes to unpredictable ruptures, restoring over 90 percent of service coverage following random node failures.
@@
  • Fairness Optimization via Structural Reconfiguration: Our analysis reveals that in crowded urban environments, the framework strategically re-optimizes spatial deployment after failure to mitigate co-channel interference, effectively surpassing its pre-failure fairness performance.
@@
The remaining sections of this paper are organized as follows. Section II reviews related work on UAV deployment and MARL for UAV swarms. Sections III and IV respectively present the system model and problem formulation. Section V details the proposed TAG-MAPPO framework. Section VI discusses the simulation setup and performance evaluation. Finally, Section VII concludes the paper.
II Related Work
@@
This section reviews existing literature across three critical domains, including UAV coverage optimization, network resilience under node failure, and the integration of graph-based attention mechanisms within multi-agent reinforcement learning (MARL).
II-A Multi-UAV Control and Coverage Optimization
@@
The deployment of UAVs for wireless coverage has been studied via heuristic, mathematical, and learning-based approaches. Early research formulated the coverage problem as a facility location task, deriving optimal 3D locations for static users via convex optimization and circle packing theory [20, 2]. To address energy constraints, trajectory optimization frameworks balanced throughput against propulsion energy consumption [31]. More recently, meta-heuristic algorithms have been applied to dynamic scenarios, such as collaborative differential evolution for path planning and energy-efficient graph algorithms for minimum-UAV coverage [27, 9]. However, these model-based methods typically assume global Channel State Information (CSI) availability and suffer from high computational latency, making them less suitable for real-time control in volatile environments.
@@
With deep reinforcement learning (DRL) advancements, decentralized control strategies have emerged as a promising alternative. Since the introduction of Multi-Agent Deep Deterministic Policy Gradient (MADDPG) [17], various DRL-based controllers have addressed adaptive 3D placement in 6G small cells [10] and joint trajectory design in space-air-ground integrated networks [16]. Beyond throughput, ensuring fairness has gained attention via adaptive deployment [13] and utility-driven route adaptation for green cooperative networks [3]. Despite these advances, most existing works rely on coordinate-based Multi-Layer Perceptrons (MLPs). This architecture struggles to capture the permutation-invariant nature of UAV swarms [29], often leading to coordination collapse when agent populations fluctuate.
II-B Network Resilience and Fault Tolerance
@@
In mission-critical scenarios, system resilience, referring to the ability to recover from sudden disruptions, is paramount. Robustness in UAV search path planning has been enhanced using DRL for disaster environments [15], while two-stage optimization has addressed collaborative post-disaster communications [34]. A recent study [11] addressed resilient topology configuration in 3D UAV networks using mixed-integer non-linear programming (MINLP) and recursive optimization to maintain connectivity. While providing theoretical bounds, such model-based approaches entail significant overhead during large-scale topological shifts.
@@
To adapt to non-stationary environments, continual learning (CL) and distributed techniques have been adopted. A resilient topology optimization framework leveraging quantum annealing [32] was introduced for rapid online switching, yet it relies on pre-defined candidate sets and specialized hardware. Building on foundational CL algorithms, a spatiotemporal-aware DRL framework [5] was introduced for cooperative coverage. In our prior work [14], a spatiotemporal CL framework based on Group-Decoupled MAPPO (G-MAPPO) enabled UAVs to adapt to varying user distributions. However, current research predominantly focuses on macro-spatial variations. Micro-temporal resilience, involving immediate self-healing after node failure without pre-computed sets, remains under-explored. This paper addresses the gap by proposing a topology-aware controller for real-time fault recovery.
II-C Attention Mechanisms and Reconfigurable MARL
@@
To overcome the limitations of fixed-input architectures, Graph Neural Networks (GNNs) and attention mechanisms have been integrated into MARL. Recent reviews emphasize that graph-based learning is uniquely suited for wireless resource management due to its inherent ability to generalize to dynamic network statuses and scale with varying numbers of nodes [7]. The effectiveness of using graph representations for resilient coordination is further supported by the topology sampling logic explored in [32]. The Transformer architecture [25] demonstrated the power of attention in sequence modeling, and it has been further shown that graph-based approaches possess strong relational inductive biases suitable for physical systems [4].
@@
In the context of UAV mobility management, a hierarchical multi-agent DRL approach was proposed in [18] to handle dynamic clustering. Although hierarchical structures can be effective, they are often complex to train. Attention mechanisms offer a more scalable solution. Algorithms such as MAAC [12] and MAGIC [22] use attention to learn dynamic communication graphs. Nevertheless, few existing works explicitly integrate relative velocity into the attention mechanism for resilient trajectory control. By embedding velocity vectors and spatial relations, the proposed TAG-MAPPO framework enables agents to perceive the evolution of the network topology. This approach enhances stability in high-mobility environments and ensures that the remaining swarm can autonomously re-optimize coverage after a catastrophic node failure.
III System Model
III-A 3D Aerial-Ground Network Architecture
III-A1 Scenario Description
@@
We consider a dynamic High-Mobility Edge Coverage scenario within 6G Non-Terrestrial Networks (NTN). As illustrated in Fig. 1, a fleet of Unmanned Aerial Vehicles (UAVs) operates as a flying small-cell tier to assist a macro Ground Base Station (GBS). This architecture is specifically designed for dynamic clusters such as vehicular platoons or emergency fleets to address severe spatiotemporal non-stationarity. The UAV swarm dynamically tracks user mobility to offload traffic and maintain reliable Line-of-Sight (LoS) connectivity, mitigating the inherent coverage limitations of terrestrial infrastructure.
III-A2 Physical Configuration and Survivability
@@
The system comprises $K_{U}$ rotary-wing UAVs and one stationary GBS, totaling $K=K_{U}+1$ serving nodes. These nodes serve $M$ ground users distributed over a square target area $\mathcal{D}\subset\mathbb{R}^{2}$. Let $\mathcal{K}_{U}=\{1,\dots,K_{U}\}$ and $\mathcal{M}=\{1,\dots,M\}$ denote the sets of mobile UAVs and users, respectively. The GBS is situated at the center $\mathbf{p}_{\mathrm{GBS}}=[x_{c},y_{c},0]^{T}$ to provide stable backhaul connectivity.
@@
System resilience is characterized by a node survival indicator $\xi_{k}(t)\in\{0,1\}$ for each UAV $k\in\mathcal{K}_{U}$. If UAV $k$ is operational at time $t$, $\xi_{k}(t)=1$; otherwise, $\xi_{k}(t)=0$ represents hardware failure or energy depletion. Since the GBS has a persistent power supply, we define $\xi_{\mathrm{GBS}}(t)=1$ for all $t$.
@@
The active subset of serving nodes at time $t$ is $\mathcal{V}(t)=\{k\in\mathcal{K}_{U}\mid\xi_{k}(t)=1\}\cup\{\mathrm{GBS}\}$. An inactive UAV ($\xi_{k}(t)=0$) ceases all locomotion and transmission functions, removing it from the coordination process and interference environment. The state of each operational UAV $k\in\mathcal{V}(t)\setminus\{\mathrm{GBS}\}$ includes its 3D coordinates $\mathbf{p}_{k}(t)=[x_{k}(t),y_{k}(t),z_{k}(t)]^{T}$ and instantaneous velocity $\mathbf{v}_{k}(t)$, subject to a maximum flight speed $\|\mathbf{v}_{k}(t)\|\leq V_{\max}^{\text{UAV}}$ and a safe altitude corridor $H_{\min},H_{\max}]$.
III-A3 Inter-UAV Coordination and Control Plane